OpenStudy (anonymous):

Please solve the following absolute value inequalitys. 1.) |4m+2|>14 2.) |4m-2|<14 (Take note: The ">, & <" have the _ (line) underneath the symbol.)

OpenStudy (anonymous):

\[|4x+2|\geq 14\] \[4x+2\geq -14\] or \[4x+2\geq 14\] solve separately \[4x+2\leq -14\] \[4x\leq -16\] \[x\leq -4\] or \[4x+2\geq 14\] \[4x\geq 12\] \[x\geq 3\]

OpenStudy (anonymous):

answer is \(x\leq -4\) or \(x\geq 3\)

OpenStudy (anonymous):

\[|4m-2|\leq14\] \[-14\leq 4m-2\leq 14\] \[-12\leq 4m\leq 16\] \[-3\leq m\leq 4\]